What is the theoretical yield of vanadium that can be produced by the reaction of 40g of V2O5 with 40g of calcium based on the following chemical equation? - V2O5 + 5Ca -----> 2V + 5CaO


1
Expert's answer
2021-05-25T07:49:49-0400

Moles of "V_2O_5= \frac{40}{181.88}=0.2199" mol


Moles of "Ca= \frac{40}{40.078}=0.9980mol"




Hence Ca is the limiting reactant


"V_2O_5 + 5Ca ---> 2V + 5CaO"


From the equation,

Mass of "V_2O_5" produced = "\frac{1}{5} ×0.9980×181.88 g/mol =36.30 g"